# Pseudodifferential damping estimates and stability of relaxation shocks

### v2

@@ Abstract.
A bottleneck in the theory of large-amplitude and multi-D viscous and relaxation shock stability is the development of nonlinear damping estimates controlling higher by lower derivatives. These have traditionally proceeded from time-evolution bounds based on Friedrichs symmetric and Kawashima or Goodman type energy estimates. Here, we propose an alternative program based on frequency-dependent pseudodifferential time-space damping estimates in the spirit of Kreiss. These are seen to be equivalent in the linear case to high-frequency spectral stability, and, just as for the constant-coefficient analysis of Kreiss, sharp in a pointwise, fixed-frequency, sense. This point of view leads to a number of simplifications and extensions using both already-existing analysis and new results developed here. In particular, we essentially resolve fully existence of damping estimates for smooth, noncharacteristic relaxation waves in the small-amplitude case for multi-D and in the large-amplitude case for 1-D, showing that both reduce to satisfaction of a standard linear-algebraic dissipativity condition at the endstates of the shock, related to high-frequency stability of constant solutions. For large-amplitude multi-D shocks, we point out a key new difficulty in the form of Airy-type turning points analogous to Kreiss’ glancing points in the constant-coefficient case, and introduce a variable-coefficient adaptation of Kreiss’ symmetrizers allowing for the first time the treatment of large-amplitude smooth noncharacteristic relaxation shocks in multi-D.
@@ 1.1. Main results
Here, we address this bottleneck with the addition of new tools, restricting for simplicity to the case of smooth waves, for which we need consider only interior structure without the added complication of boundary/transmission conditions. We first propose in place of (1.2) a pseudodifferential analog (2.9) that can be shown to be linearly equivalent to (1.3). See Proposition 2.1. Next, we show that existence of a multi-dimensional Kreiss symmetrizer implies a time-integrated version of the classical nonlinear damping estimate, and discuss relations to exponential dichotomies and WKB expansion. See Propositions 3.3 and 3.7.
@@
In the large-amplitude multi-D case, we point out the new issue of Airy-type turning points, and an associated distinction between exponential dichotomies and symmetrizers, namely, that existence of dichotomies is strictly stronger than existence of symmetrizers, so that symmetrizer construction is the tool of more general application. See Propositions 5.2 and 5.5. Combining exponential dichotomies with Kreiss-type constant-coefficient symmetrizers, we carry out a symmetrizer construction in a generic case involving single Airy-type turning points. See Theorem 6.2. This makes possible for the first time the treatment of nonlinear stability of large-amplitude multi-D relaxation shocks. In particular, our results apply to arbitrary-amplitude smooth hydraulic shock waves of the Saint Venant equations, as studied in [54]. We show in Section 7 that the time-integrated nonlinear damping estimate resulting from our pseudodifferential damping condition serves equally well as the classical one to close a nonlinear iteration and yield a nonlinear stability result. See Theorem 7.2. We discuss applications and larger perspectives in Sections 8-9.
@@ 2. Pseudodifferential damping and high-frequency stability
Consider following the approach of [33, 58] a nearby wave $w(x+\psi(x,t),t)=\bar{w}(x)+v(x,t)$, where $v$ is a nonlinear residual and $\psi$ is a possible small phase modulation depending on $v$ and $\bar{w}$- necessary in 1-D, but typically not needed in multi-D [58]. Then, the nonlinear perturbation equations governing $v$ take the form
@@
where $A_{j}=(df_{j}/dw)(\bar{w}+v)$, $Ev=-(dr/dw)(\bar{w})v+(d^{2}f_{1}/dw^{2})(v,\partial_{x_{1}}\bar{w})$, and $f$ is a quadratic order remainder term in $v$ and derivatives of $\psi$ multiplied by factor $\partial_{x_{1}}\bar{w}$, depending also on $\bar{w}$ and its derivatives, but not derivatives of $v$ and $\psi$.
@@ 2.1.1. Time-integrated version
(2.3) $\|v(T)\|_{H^{s}_{\alpha}}^{2}\leq Ce^{-\eta T}\|v(0)\|_{H^{s}_{\alpha}}^{2}+C\int_{0}^{T}e^{-\eta(T-s)}\big(\|v(t)\|_{L^{2}_{\alpha}}^{2}+\|f(t)\|_{H^{s}_{\alpha}}^{2}\big)\,dt$
@@ 2.2. Paralinearization
To remove extraneous symmetry assumptions and connect directly to the linearized resolvent equation, we perform a pseudodifferential analysis, replacing $\partial_{x_{j}}$, $j=2,\dot{,}d$ and $\partial_{t}$ by their symbols $i\eta_{j}$ and $i\tau$. See [41, Appendix B] for a description of the relevant tools, phrased in the paradifferential calculus of Bony; different from the setting of [41], there is no additional parameter here such as viscosity or relaxation constant, simplifying slightly the situation. This has the added advantage, in the case of planar traveling-wave profiles containing discontinuous subshocks, of putting things in a common framework with the finite-time stability analysis of discontinuous shock problems pioneered in [23, 31, 32, 37], where frequency-dependent (in time) estimates are seen to be necessary for multi-D stability estimates.
@@ 2.3. Pseudodifferential damping
In the short-time discontinuous theory of [23, 31, 32, 37], the goal is, by frequency-dependent type symmetric energy estimates (see Method of symmetrizers below) to obtain uniform “semigroup type” estimates
@@
for (2.4), and analogous estimates for (2.5) involving also $g$. Parseval’s identity gives then
@@ 3.1. Symmetrizers
We begin by repeating essentially verbatim the generalization of Kreiss’ method to variable-coefficient problems in [40, §2.1], there stated for the discontinuous case. In the simpler smooth case, the spatial domain is extended to $(-\infty,\infty)$ and (3.4) dropped. Recall now the essence of the “method of symmetrizers” as it applies to general boundary value problems
@@ 3.2.1. The quantitative estimates of Lannes
For further details, see the commutator estimates of [28, Thm. 7, p. 529], the quantity $H^{t_{0}}_{ref}$ defined in [28, Eq. (4.7), p. 519], and the quantities $n$ and $N$ on which $H^{t_{0}}_{ref}$ depends, defined in [28, Eqs. (2.2) and (2.4), pp. 502 and 503]; see also [28, Thms. 1 and 2, pp. 514 and 515] and [28, Cor. 30, p. 515] bounding the action of operators in Sobolev norms in terms of Sobolev norms of their symbols. For the main, $H^{s}$ (rather than $H^{s-1}$ of gradient) bounds, see also the alternative, more streamlined treatment of Sroczinski in [50, §2], focused entirely on Sobolev-based bounds. For general discussions of pseudo- and paradifferential calculus, see [3, Appendix C] and [39].
@@ 3.2.2. Reduction to perturbed linear case
3.3. The small-amplitude case
In the small-amplitude case that the entire profile $\bar{w}(x_{1})$ lies close to a constant (necessarily equlibrium) solution $w(x_{1})\equiv w_{0}$, with $\|\bar{w}-w_{0}\|_{W^{1,\infty}}$ sufficiently small, we can make a much stronger statement. For, in this case we may paralinearize in all coordinates $x_{1},\dots,x_{d},t$ to reduce to an entirely symbolic computation, since commutator terms, controlled by Lipshitz norm are then absorbable errors in the estimate.
@@ 3.3.1. The symmetric case
This may be compared with the standard ($G$ rather than $\tilde{G}$) symmetrizer construction of [42, Appendix E] for hyperbolic-parabolic, or “viscous” conservation laws
@@ 3.4. Partial converse and connection to exponential dichotomies
Proposition 3.7.
Existence of pointwise exponential dichotomies is necessary for the high-frequency resolvent bound (1.3), which is necessary for existence of symmetrizers. On the other hand, existence of uniformly smooth and bounded global exponential dichotomies is sufficient for existence of symmetrizers.
@@
(3.21) $\langle z_{+}(x),Q_{+}(x)z_{+}(x)\rangle$ $=\int_{x}^{+\infty}\langle z_{+}(x),{\mathcal{S}}_{+}^{T}(y-x){\mathcal{S}}_{+}(y-x)z_{+}(x)\mathrm{range}dy$
@@ 4. The large-amplitude 1-D noncharacteristic hyperbolic case
Theorem 4.1.
In one space dimension, assuming (A1)–(A3) near endstates $w_{\pm}$ and (A1)–(A2), (A3’) near points on the orbit of $\bar{w}(\cdot)$, high-frequency resolvent estimate (1.3), pseudodifferential damping (2.9), existence of smooth symmetrizers, and existence of uniform smooth exponential dichotomies are all equivalent, and, moreover, are equivalent to
@@
we may thus appeal to the tracking/reduction lemma of [36, §3.2] to conclude the existence of an ${\rm Id}+O(\varepsilon)$ coordinate change converting (4.11) to exact diagonal form

### v3

@@ Abstract.
A bottleneck in the theory of large-amplitude and multi-D viscous and relaxation shock stability is the development of nonlinear damping estimates controlling higher by lower derivatives. These have traditionally proceeded from time-evolution bounds based on Friedrichs symmetric and Kawashima or Goodman type energy estimates. Here, we propose an alternative program based on frequency-dependent pseudodifferential time-space damping estimates in the spirit of Kreiss. These are seen to be equivalent in the linear case to high-frequency spectral stability, and, just as for the constant-coefficient analysis of Kreiss, sharp in a pointwise, fixed-frequency, sense. This point of view leads to a number of simplifications and extensions using both already-existing analysis and new results developed here. In particular, we essentially resolve fully existence of damping estimates for smooth, noncharacteristic relaxation waves in the small-amplitude case for multi-D and in the large-amplitude case for 1-D, showing that both reduce to satisfaction of a standard linear-algebraic dissipativity condition at the endstates of the shock, related to high-frequency stability of constant solutions. For large-amplitude multi-D shocks, we point out a key new difficulty in the form of Airy-type turning points analogous to Kreiss’ glancing points in the constant-coefficient case, and introduce a variable-coefficient adaptation of Kreiss’ symmetrizers allowing for the first time the treatment of large-amplitude noncharacteristic relaxation shocks in multi-D.
@@ 1.1. Main results
Here, we address this bottleneck with the addition of new tools, for simplicity restricting mainly to the case of smooth waves, for which we need consider only interior structure without the added complication of boundary/transmission conditions. We first propose in place of (1.2) a pseudodifferential analog (2.9) that can be shown to be linearly equivalent to (1.3). See Proposition 2.1. Next, we show that existence of a multi-dimensional Kreiss symmetrizer implies a time-integrated version of the classical nonlinear damping estimate, and discuss relations to exponential dichotomies and WKB expansion. See Propositions 3.3 and 3.7.
@@
In the large-amplitude multi-D case, we point out the new issue of Airy-type turning points, and an associated distinction between exponential dichotomies and symmetrizers, namely, that existence of dichotomies is strictly stronger than existence of symmetrizers, so that symmetrizer construction is the tool of more general application. See Propositions 5.2 and 5.5. Combining exponential dichotomies with Kreiss-type constant-coefficient symmetrizers, we carry out a symmetrizer construction in a generic case involving single Airy-type turning points. See Theorem 6.2. This makes possible for the first time the treatment of nonlinear stability of large-amplitude multi-D relaxation shocks. In particular, our results apply to arbitrary-amplitude smooth hydraulic shock waves of the Saint Venant equations, as studied in [55]. We show in Section 7 that the time-integrated nonlinear damping estimate resulting from our pseudodifferential damping condition serves equally well as the classical one to close a nonlinear iteration and yield a nonlinear stability result. See Theorem 7.2.
Along the way, we discuss in Sections 2.5.1, 3.4.1, 4.1, 6.1, and 7.3 the changes needed for extension to the discontinuous case, in particular sketching the treatment in 1- and multi-D of arbitrary-amplitude shock profiles with a single subshock discontinuity, of the types arising for discontinuous hydraulic shock waves of the Saint Venant equations [54, 11, 55]. See Proposition 3.9, Corollary 4.3, and the treatment in Theorem 6.4 of a multi-d turning point case scenario noted by Erpenbeck [8, 27, 61] in the case of detonations that as shown in [55] is relevant for discontinuous hydraulic shock profiles of the Saint Venant equations of inclined shallow-water flow. See Theorem 6.4.
Finally, we discuss applications and larger perspectives in Sections 8-9, in particular pointing out as an important open problem the complete treatment of nonlinear stability of large-amplitude discontinuous profiles in multi-D, building on the nonlinear damping estimates established here.
@@ 2. Pseudodifferential damping and high-frequency stability
Consider following the approach of [34, 59] a nearby wave $w(x,t)$ and perturbation $v$ defined by
$w(x_{1}+\psi(x_{2},\dots,x_{d},x_{2},t),x_{2},\dots,x_{d},,t)=\bar{w}(x)+v(x,t),$
where $v$ is a nonlinear residual and $\psi$ is a possible small phase modulation depending on $v$ and $\bar{w}$.
Then, the nonlinear perturbation equations governing $v$ take the form
@@
where
$A_{1}=(df_{1}/dw)(\bar{w}+v)+\partial_{t}\psi\,{\rm Id}+\sum_{j=2}^{d}\partial_{x_{j}}\psi\,(df_{j}/dw)(\bar{w}+v),$
$A_{j}=(df_{j}/dw)(\bar{w}+v)$ for $j=2,\dots,d$, $Ev=-(dr/dw)(\bar{w})v+(d^{2}f_{1}/dw^{2})(v,\partial_{x_{1}}\bar{w})$, and $f$ is a quadratic order remainder term in $v$ and derivatives of $\psi$ multiplied by factor $\partial_{x_{1}}\bar{w}$, depending also on $\bar{w}$ and its derivatives, but not on $\psi$ or derivatives of $v$.
The incorporation of a phase-modulation $\psi$ is necessary in 1-D or discontinuous cases, but typically not needed for smooth waves in multi-D [59]. For a standard multi-D analysis of smooth waves– our main object here– the reader may consider for simplicity $\psi(x,t)\equiv 0$.
@@ 2.1.1. Time-integrated version
(2.3) $\|v(T)\|_{H^{s}_{\alpha}}^{2}\leq Ce^{-\eta T}\|v(0)\|_{H^{s}_{\alpha}}^{2}+C\int_{0}^{T}e^{-\eta(T-t)}\big(\|v(t)\|_{L^{2}_{\alpha}}^{2}+\|f(t)\|_{H^{s}_{\alpha}}^{2}\big)\,dt$
@@ 2.2. Paralinearization
To remove extraneous symmetry assumptions and connect directly to the linearized resolvent equation, we perform a pseudodifferential analysis, replacing $\partial_{x_{j}}$, $j=2,\dots,d$ and $\partial_{t}$ by their symbols $i\eta_{j}$ and $i\tau$. See [42, Appendix B] for a description of the relevant tools, phrased in the paradifferential calculus of Bony; different from the setting of [42], there is no additional parameter here such as viscosity or relaxation constant, simplifying slightly the situation. This has the added advantage, in the case of planar traveling-wave profiles containing discontinuous subshocks, of putting things in a common framework with the finite-time stability analysis of discontinuous shock problems pioneered in [24, 32, 33, 38], where frequency-dependent (in time) estimates are seen to be necessary for multi-D stability estimates.
@@ 2.3. Pseudodifferential damping
In the short-time theory of [24, 32, 33, 38], the goal is, by frequency-dependent type symmetric energy estimates (see Method of symmetrizers below) to obtain uniform “semigroup type” estimates
@@
for (2.4), and analogous estimates (2.12) for (2.5) involving also $g$. Parseval’s identity gives then
@@ 2.5. Linear equivalence
2.5.1. The discontinuous case
For the case (2.5) of a single discontinuity, the analog of (2.6) is
(2.12) $\|v\|_{\hat{H}^{s}}\leq\frac{C(\|f\|_{\hat{H}^{s}}+|g|_{\hat{H}^{s}})}{\Re\lambda-\gamma_{*}},$
where $\|f\|_{\hat{H}^{s}}:=\|f\|_{H^{s}}+(1+|\xi|)^{s}\|f\|_{L^{2}}$ and $|g|_{\hat{H}^{s}}:=(1+|\xi|)^{s}|g|_{L^{2}}$, and the analog of (2.9) is
(2.13) $\|v\|_{\hat{H}^{s}}\leq\frac{C(\|f\|_{\hat{H}^{s}}+|g|_{\hat{H}^{s}}+\|v\|_{L^{2}})}{\Re\lambda-\gamma_{*}},$
for some $\gamma_{*}<0$. Evidently, the same (tautological) argument as in the proof of Proposition 2.1 gives linear equivalence of these two conditions as well, as it makes no use of boundary conditions.
@@ 3.1. Symmetrizers
We begin by repeating essentially verbatim the generalization of Kreiss’ method to variable-coefficient problems in [41, §2.1], there stated for the discontinuous case. In the simpler smooth case, the spatial domain is extended to $(-\infty,\infty)$ and (3.4) dropped. Recall now the essence of the “method of symmetrizers” as it applies to general boundary value problems
@@ 3.2.1. The quantitative estimates of Lannes
For further details, see the commutator estimates of [29, Thm. 7, p. 529], the quantity $H^{t_{0}}_{ref}$ defined in [29, Eq. (4.7), p. 519], and the quantities $n$ and $N$ on which $H^{t_{0}}_{ref}$ depends, defined in [29, Eqs. (2.2) and (2.4), pp. 502 and 503]; see also [29, Thms. 1 and 2, pp. 514 and 515] and [29, Cor. 30, p. 515] bounding the action of operators in Sobolev norms in terms of Sobolev norms of their symbols. For the main, $H^{s}$ (rather than $H^{s-1}$ of gradient) bounds, see also the alternative, more streamlined treatment of Sroczinski in [51, §2], focused entirely on Sobolev-based bounds. For general discussions of pseudo- and paradifferential calculus, see [3, Appendix C] and [40].
@@ 3.2.2. Reduction to perturbed linear case
3.3. The smooth small-amplitude case
In the smooth small-amplitude case that the entire profile $\bar{w}(x_{1})$ lies close to a constant (necessarily equlibrium) solution $w(x_{1})\equiv w_{0}$, with $\|\bar{w}-w_{0}\|_{W^{1,\infty}}$ sufficiently small, we can make a much stronger statement. For, in this case we may paralinearize in all coordinates $x_{1},\dots,x_{d},t$ to reduce to an entirely symbolic computation, since commutator terms, controlled by Lipshitz norm are then absorbable errors in the estimate.
@@ 3.3.1. The symmetric case
This may be compared with the standard ($G$ rather than $\tilde{G}$) symmetrizer construction of [43, Appendix E] for hyperbolic-parabolic, or “viscous” conservation laws
@@ 3.4. Partial converse and connection to exponential dichotomies
Proposition 3.7 (Smooth case).
For smooth profiles, existence of pointwise exponential dichotomies on $(-\infty,\infty)$ is necessary for the high-frequency resolvent bound (1.3), which is necessary for existence of symmetrizers. On the other hand, existence of uniformly smooth and bounded global exponential dichotomies is sufficient for existence of symmetrizers.
@@
(3.21) $\langle z_{+}(x),Q_{+}(x)z_{+}(x)\rangle$ $=\int_{x}^{+\infty}\langle z_{+}(x),{\mathcal{S}}_{+}^{T}(y-x){\mathcal{S}}_{+}(y-x)z_{+}(x)\rangle dy$
@@
3.4.1. The discontinuous case
Similarly as in the analogous case of a spectral gap in the constant-coefficient case, it is straightforward to extend the “interior symmetrizer construction” just performed to one achieving also boundary coercivity, (3.4), at least in the case of a profile containing a single “subshock” discontinuity, without loss of generality at the origin, $x_{1}=0$.
Recall [24, 32, 33, 38], even for well-posedness/short-time existence of the linearized equations, the subshock must be of Lax type, satisfying the linearized Rankine-Hugoniot conditions
(3.26) $\psi[\lambda F(\bar{w})+\sum_{j=2}^{d}\xi_{j}F_{j}(\bar{w})]+[A_{1}v]=0,$
where $\cdot$] denotes jump across $x=0$ and $\psi$ the infinitesimal shock location, from which, projecting onto directions orthogonal to $\lambda F(\bar{w})+\sum_{j=2}^{d}\xi_{j}F_{j}(\bar{w})]$ to eliminate $\psi$, one obtains a rank-(n1)(n-1) frequency-dependent boundary condition
$\Gamma(\lambda,\xi_{2},\dots,\xi_{d})v=0$
on the traces $(v(0^{+}),v(0^{-}))$.
By Palmer’s theorem [47], we obtain existence of exponential dichotomies on $(-\infty,0)$ and $(0,+\infty)$. The linear high-frequency stability estimate (2.12), specialized to the homogeneous case $f=0$, $g=0$, then gives $v\equiv 0$ for $H^{s}$ solutions $v$ of (2.5), from which we obtain the (variable-coefficient) Kreiss-Lopatinsky condition
(KL) $\hbox{\rm$\Gamma$ is full rank on $(\Sigma_{+}(0^{-}),\Sigma_{-}(0^{+}))$},$
where $\Sigma_{\pm}:={\rm Range}P_{\pm}$, with $P_{\pm}$ the projectors onto exponentially growing (decaying) modes in forward $x_{1}$.
Consulting to the block-diagonaling coordinates of (3.17)-(3.19), we have that (KL) is equivalent to $\Gamma$ full rank on the image under $T$ of the coordinate blocks associated with $Q_{+}$ for $x<0$ and $Q_{-}$ for $x>0$. Thus, redefining (3.23) as
(3.27) $S(x)$ $=\Big(T^{-1,T}\begin{pmatrix}-Q_{+}&0\\ 0&CQ_{-}\end{pmatrix}T\Big(x),\quad x<0,$
$S(x)$ $=\Big(T^{-1,T}\begin{pmatrix}CQ_{+}&0\\ 0&-Q_{-}\end{pmatrix}T\Big(x),\quad x>0$
with $C>0$ sufficiently large, we obtain (3.4), exactly as in the treatment by Kreiss of the constant coefficient case with spectral gap. (See first part of proof of Main Theorem 4 [24, top, p. 292].)
Combining, we have the following discontinuous version of Proposition 3.7.
Proposition 3.9 (Discontinuous case).
For profiles with a single discontinuity at $x=0$, existence of pointwise exponential dichotomies on $(-\infty,0)$ and $(0,+\infty)$ together with the Kreiss-Lopatinsky condition are necessary for the high-frequency resolvent bound (1.3), which is necessary for existence of symmetrizers. On the other hand, existence of (globally) uniformly smooth and bounded exponential dichotomies on on $(-\infty,0)$ and $(0,+\infty)$ together with the Kreiss-Lopatinsky condition is sufficient for existence of symmetrizers.
3.4.2. The case of multiple discontinuities
We mention in passing the situation of a steady profile with multiple Lax-type subshock discontinuities, which cannot happen in the constant-coefficient case, but can happen for relaxation profiles [20, 11]. For instance, let us consider the case of two shocks, at $x_{1}$ and $x_{2}>x_{1}$. In this case, the Kreiss-Lopatinski condition (KL) is replaced by a more complicated rank condition involving the reduced Rankine-Hugoniot conditions $\Gamma_{j}$ for both shocks, $j=1,2$, and the values of $v$ on the growing manifold to the left of $x_{1}$, the decaying manifold to the right of $x_{2}$. This requires a more detailed computation than the one for single shocks, involving also dynamics on the intermediate zone between the shocks. We return to this issue in the more concrete setting of Section 4.2, in which these intermediate zone dynamics may be diagonalized.
@@ 4. The large-amplitude 1-D noncharacteristic hyperbolic case
Theorem 4.1 (Smooth case).
For smooth profiles in 1D, assuming (A1)–(A3) near endstates $w_{\pm}$ and (A1)–(A2), (A3’) near points on the orbit of $\bar{w}(\cdot)$, high-frequency resolvent estimate (1.3), pseudodifferential damping (2.9), existence of smooth symmetrizers, and existence of uniform smooth exponential dichotomies are all equivalent, and, moreover, are equivalent to
@@
we may thus appeal to the tracking/reduction lemma of [37, §3.2] to conclude the existence of an ${\rm Id}+O(\varepsilon)$ coordinate change converting (4.11) to exact diagonal form
@@
4.1. The discontinuous case
For a profile with a single discontinuity at $x=0$, we obtain by essentially the same construction of exponential dichotomies as in the proof of Proposition (4.1), now on half-lines $(-\infty,0)$ and $(0,+\infty)$, together with Proposition 3.9, an equally decisive result.
Corollary 4.3 (Discontinuous case).
For profiles in 1D with a single discontinuity (without loss of generality at $x=0$), assuming (A1)–(A3) near endstates $w_{\pm}$ and (A1)–(A2), (A3’) near points on the orbit of $\bar{w}(\cdot)$, high-frequency resolvent estimate (1.3), pseudodifferential damping (2.9), and existence of smooth symmetrizers are all equivalent, and, moreover are equivalent to high-frequency dissipativity conditions (4.3) (resp. (4.4) in the case of multiple eigenvalues) at the endstates $w_{\pm}$ together with the Kreiss-Lopatinsky condition (KL).
4.2. Multiple subshocks
In the 1-D strictly hyperbolic case we are considering, for which we may reduce as above to diagonal form, one may for low-dimensional systems carry out a classical damping estimate for solutions with two subshocks by the methods used in [50], whenever there holds the high-frequency resolvent estimate (1.3). For $2\times 2$ systems with two subshocks, the same technique can be adapted to yield symmetrizers, and therefore pseudodifferential damping, whenever there holds (1.3). We demonstrate here the essential part of the argument by treating the case of a bounded interval $x\in[0,L]$ and diagonal system
(4.14) $u^{+}_{t}+a_{+}u^{+}_{x}=-b^{+}u^{+},$
$u^{-}_{t}+a_{-}u^{-}_{x}=-b^{-}u^{-},$
$a^{\pm}\gtrless 0$, $b^{\pm}>0$, with boundary conditions
(4.15) $u^{-}(0)$ $=c^{-}u^{+}(0),$
$u^{+}(L)$ $=c^{+}u^{-}(L).$
This yields eigenvalue equations
(4.16) $(u^{\pm})_{x}=g^{\pm}u^{\pm},\qquad g^{\pm}=-(a^{+})^{-1}(\lambda+b^{+}),$
with boundary conditions (4.15), for which (1.3) implies the Evans-Lopatinsky condition of nonexistence of solutions for $\Re\lambda\geq 0$ and $|\lambda|$ sufficiently large. By explicit exponentiation of the eigenvalue problem (4.16) and examination of (4.15), one finds as in [50] that this amounts to
(4.17) $|c^{+}c^{-}|<|v^{+}(L)|/|v^{-}(L)|,$
where $v^{\pm}$ are solutions of (4.16) with $\lambda=0$, initialized as $v^{\pm}(0)=1$.
Alternative symmetrizer construction. In this diagonal setting, we may construct a symmetrizer in a more precise and simple way, following the idea of [50], as $S={\rm diag}\{s^{+},s^{-}\}$, with $s^{\pm}$ defined by
(4.18) $2\Re s^{\pm}g^{\pm}+s^{\pm}_{x}=\varepsilon,$
giving in the zero-interior dissipation limit $\varepsilon\to 0$ solutons of
$s^{\pm}_{x}=-2\Re s^{\pm}g^{\pm}=-2(b^{\pm}/a^{\pm})\,s^{\pm},$
hence
(4.19) $s^{\pm}=|d^{\pm}/v^{\pm}|^{2},$
with $d^{\pm}$ constant, without loss of generality $^{-}=1, d+=d$$^{+}=d$.
It remains to show strict boundary dissipativity for this $\varepsilon\to 0$ limit, from which we obtain boundary dissipativity by continuity also for $\varepsilon>0$ sufficiently small, completing the proof. It is straightforward to see that boundary dissipativity (3.4) is equivalent to maximum dissipativity: $S(0)$ positive definite on $\hbox{\rm Ker}\Gamma(0)$ and $S(L)$ negative definite on $\hbox{\rm Ker}\Gamma(L)$, where
$\Gamma(0)=(-c^{-},1)^{T},\quad\Gamma(L)=(1,-c^{+})^{T},$
or
$\hbox{\rm Ker}\Gamma(0)=(1,c^{-})^{T},\quad\Gamma(L)=(c^{+},1)^{T},$
Taking the neutral limit
(4.20) $\Gamma(0)^{T}S(0)\Gamma(0)=0,$
we find by a similar continuity argument that it is sufficient that $\Gamma(L)^{T}S(L)\Gamma(L)<0$, or
(4.21) $|c_{+}d/v^{+}|^{2}-|1/v^{-}|^{2}<0.$
Equality (4.20) gives $d=c^{-}$, whence (4.21) becomes $|c^{+}c_{-}|<|v_{+}/v_{-}|$, in agreement with (4.17).
This computation readily extends to the two-subshock case on the whole line. More important for us, it will give a useful model for a key multi-D computation in Section 6.1.1 later on.
Remark 4.4.
For the above symmetrizer construction, it is necessary that both the length of the domain and the real part of $g^{\pm}$ coefficients be bounded, as otherwise $\max\{1/|v^{\pm}|\}$ would grow without bound, even in the constant-coefficient case. By contrast, the “Lyapunov-type” symmetrizers constructed by the method of Section 3.4 remain always bounded, in the constant-coefficient case reducing to constants. The extension to a pair of subshocks at $x_{1}<x_{2}$ within a profile defined on the whole real line may be accomplished by combining the present construction on the intermediate zone $x_{1},x_{2}]$ between the shocks with Lyapunov-type constructions on $(-\infty,x_{1})$ and $x_{2},+\infty)$, with a matching condition at the boundary points $x_{1},x_{2}$. Treatment of three or more shocks is an interesting linear algebraic problem.